[Hallc_running] [New Logentry] RC Daily Update

cyero at jlab.org cyero at jlab.org
Wed Sep 9 09:30:04 EDT 2020

Logentry Text:
<b>Accelerator Status:</b>
  <li>Sep 08/09 Swing/Owl (yesterday, today):
      <li>Accelerator beam studies (~16:30 - 19:30)</li>
      <li>Beam restoration to Hall C ~20:50 (took ~1 hr longer than expected)</li>
      <li>2L23-5 Solenoid Fault Downtime (1:03 - 3:20)  <a href="https://logbooks.jlab.org/entry/3844293">2L23-5_Downtime</a> </li>									    
  <li>Sep 09 Day (today):
      <li>MCC 7:45 meeting updates:
	  <li> Ops will do "Beam Studies" next Tuesday Day Shift (Sep 15).</li>
	  <li> In response to the Halls B/C/D beam energy (HALLC:p for us) dropping by a few MeV (~3 MeV for C) and because of the impact on beam polarization the Halls B/C/D (drop by few percent), it was
	    decided to go back to the Injector (INJ) configuration as of 02:00 Sep 08 Owl (yesterday). This work will be done this morning at the INJ (~10:00 - 11:00) for which beam to all Halls is expected to be down
	  during said time. See PD (Jay) detailed log-entry:  <a href="https://logbooks.jlab.org/entry/3843977">HALLC:p</a> </li>

<b>General Hall C Updates:</b>
  <li> Sep 08/09 Swing/Owl (yesterday, today):
      <li>NOTE: Yesterday, B. Sawatzky noticed abnormaly high rates when beam to Hall C was OFF (during d2n temperature tests) while beam to Hall A was ON. This lead to the
	suspicion of bleed-through beam from A to C <a href="https://logbooks.jlab.org/entry/3844073">HallA_bleedthrough</a>. Will have to watch for this during future target work that requires no beam in Hall.
	Follow-up entry by William (Bill) Henry shows exactly when suspected bleed-through started (~15:36, yesterday) <a href="https://logbooks.jlab.org/entry/3844179">bleedthrough_rates</a> </li>
      <li> Junhao Chen did d2n temperature tests (~15:30 - 16:30) with Target: "Pickup Coil" inserted. The results are posted at: <a href="https://logbooks.jlab.org/entry/3844089">d2n_temp_test</a></li>
      <li> At ~16:24, field was rotated from +90 to +270 deg. <a href="https://logbooks.jlab.org/entry/3844064">270_deg_rotation</a> </li>
      <li> At ~16:36, NMR mesurement was done at +270 deg and target polarization was determined to be 33.5 %.  <a href="https://logbooks.jlab.org/entry/3844050">270_deg_targetNMR</a> </li>
      <li> Beam was restored at ~20:50 and beam profile checks were done.  <a href="https://logbooks.jlab.org/entry/3844138">IHA3H07A</a>,  <a href="https://logbooks.jlab.org/entry/3844142">IHA3H07B</a> </li>
      <li> At ~21:43, NMR measurement of Pol He3 (+270 deg) was determined to be 36.7 % polarization (still too low). <a href="https://logbooks.jlab.org/entry/3844164">NMR_meas</a> <br>
	It was decided to change to reference cell target (scraping checks: <a href="https://logbooks.jlab.org/entry/3844195">ref_cell_scraping_test</a>) and take reference cell (N2, 135 psig)
	to allow for polarization to increase closer to ~40 %.</li>      
      <li>At ~05:17, NMR measureement reported 37.6 % polarization. It was decided to start (05:20) d2n Pol He3 (+270 deg) production at HMS Kin-C and SHMS Kin-Z at IHWP = OUT configuration.</li>
      <li> Beam Current History during Sep 08/09 Swing/Owl:  <a href="https://logbooks.jlab.org/entry/3844281">Sep08_Swing</a>,  <a href="https://logbooks.jlab.org/entry/3844442">Sep09_Owl</a> </li>
  <li> Sep 09 Day (today):
      <li>Continue d2n Pol He3 (+270 deg) production at HMS Kin-C and SHMS Kin-Z at IHWP = OUT configuration. </li>
      <li>Beam is expected to be down for INJ work this morning at (~ 10:00 - 11:00).</li>

<b>Recent NMR/pNMR Measurements:</b><br>
<li>Sep 08 (Day/Swing):</li>
  <li> NMR (+90 deg):  09:24 | 40.7 %  <a href="https://logbooks.jlab.org/entry/3843811">https://logbooks.jlab.org/entry/3843811</a></li>
  <li> NMR (+90 deg):  15:40 | 39.0 %  <a href="https://logbooks.jlab.org/entry/3844047">https://logbooks.jlab.org/entry/3844047</a></li>
  <li> pNMR (+90 deg): 15:39 | 40.5 %  <a href="https://logbooks.jlab.org/entry/3844047">https://logbooks.jlab.org/entry/3844047</a></li>
  <li> NMR (+270 deg): 16:36 | 33.5 %  <a href="https://logbooks.jlab.org/entry/3844050">https://logbooks.jlab.org/entry/3844050</a></li>
  <li> NMR (+270 deg): 21:43 | 36.7 %  <a href="https://logbooks.jlab.org/entry/3844164">https://logbooks.jlab.org/entry/3844164</a></li>  

<li>Sep 09 (Owl):</li>
  <li> NMR (+270 deg): 02:21 | 37.7 %  <a href="https://logbooks.jlab.org/entry/3844308">https://logbooks.jlab.org/entry/3844308</a></li>
  <li> pNMR (+270 deg):02:21 | 31.0 %  <a href="https://logbooks.jlab.org/entry/3844308">https://logbooks.jlab.org/entry/3844308</a></li>
  <li> NMR (+270 deg): 05:17 | 37.6 %  <a href="https://logbooks.jlab.org/entry/3844366">https://logbooks.jlab.org/entry/3844366</a></li>

<b>Recent IHWP Changes:</b><br>
<li>Sep 03 (Swing, as of 19:58) | IHWP = IN </li>
<li>Sep 04 (Owl, as of 04:00)   | IHWP = OUT <a href="https://logbooks.jlab.org/entry/3840776">https://logbooks.jlab.org/entry/3840776</a> </li>
<li>Sep 04 (Day, as of 12:01)   | IHWP = IN <a href="https://logbooks.jlab.org/entry/3841011">https://logbooks.jlab.org/entry/3841011</a> </li>
<li>Sep 05 (Day, as of 14:55)   | IHWP = OUT <a href="https://logbooks.jlab.org/entry/3841696">https://logbooks.jlab.org/entry/3841696</a> </li>
<li>Sep 06 (Owl, as of 07:00)   | IHWP = IN <a href="https://logbooks.jlab.org/entry/3842175">https://logbooks.jlab.org/entry/3842175</a> </li>
<li>Sep 06 (Swing, as of 22:04) | IHWP = OUT <a href="https://logbooks.jlab.org/entry/3842667">https://logbooks.jlab.org/entry/3842667</a> </li>
<li>Sep 07 (Day, as of 12:11)   | IHWP = IN <a href="https://logbooks.jlab.org/entry/3843209">https://logbooks.jlab.org/entry/3843209</a> </li>
<li>Sep 08 (Owl, as of 05:05)   | IHWP = OUT <a href="https://logbooks.jlab.org/entry/3843688">https://logbooks.jlab.org/entry/3843688</a> </li>

Bluejeans remote connection information:  <br>
To join the Meeting: <br>
 <a href="https://bluejeans.com/861439920">https://bluejeans.com/861439920</a> <br>
To join via Room System: <br>
Video Conferencing System: bjn.vc -or- <br>
Meeting ID : 861439920 <br>
To join via phone : <br>
1) Dial: +1.408.740.7256 (US (San Jose)) <br>
+1.888.240.2560 (US Toll Free)) <br>
+1.408.317.9253 (US (Primary, San Jose)) <br>
(see all numbers -  <a href="http://bluejeans.com/numbers">http://bluejeans.com/numbers</a> ) <br>
2) Enter Conference ID : 861439920 <br>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3844471
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200909/9b530351/attachment.html>

More information about the Hallc_running mailing list